OpenStudy (anonymous):

Solve for x. 7 + 6(x + 2) = 43 A. x = 4 B. x = 5 C. x = 3 D. x = 8

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

O hai photon

OpenStudy (anonymous):

I know and don't know.

OpenStudy (photon336):

\[A(B+C) = A*b + A*C\]

OpenStudy (anonymous):

How is that? there is only on A

OpenStudy (photon336):

we need to take care of that 6(x+2) first 6x+12+7 = 43

OpenStudy (anonymous):

1) subtract 7 b) divide by 6 iii) subtract 2

OpenStudy (anonymous):

those three steps will get your answer

OpenStudy (anonymous):

What am I distributing or something?

OpenStudy (photon336):

7+6(x+2) = 43 6(x+2) = 6x+12 we're distribution the 6 in the ( )

OpenStudy (anonymous):

I saw that. How would I do it his way? :P

OpenStudy (anonymous):

XD

OpenStudy (photon336):

let me show you 7+6(x+2) = 43 -7 -7 step #1 subtract 7 6(x+2) = 36 6 6 step #2 divide by 6 x+2 = 6 step #3 subtract 2 x+2 = 6 -2 -2 x= 4

OpenStudy (photon336):

the way I showed you 7+6(x+2) = 43 7+6x+12 = 43 6x+19 = 43 -19 -19 6x = 24 6 6 x = 4

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

I can remember his way XD

OpenStudy (photon336):

you see the deal here right? two different ways of approaching the same problem and same answer.

OpenStudy (anonymous):

true

OpenStudy (photon336):

His way of doing it though is a little easier whatever works for you

pooja195 (pooja195):

@AnimeWover you you understand or do you want more help?

OpenStudy (anonymous):

I understand

pooja195 (pooja195):

Alright. Just makin sure.

OpenStudy (anonymous):

XD

OpenStudy (mathmale):

Here's an alternative approach: Subtract 7 from both sides of the given equation: 7 + 6(x + 2) = 43 Then divide the result by 6. Then what?

OpenStudy (anonymous):

um. That was just stated